OpenStudy (anonymous):

What is the slope of the line passing through the points (0, 4) and (−8, −1)?

OpenStudy (bohotness):

m = (y₂ - y₁) / (x₂ - x₁) m = (-1 - 4) / ( -8 - 0) m = -5 / -8 m = 5/8 <----- https://s.yimg.com/hd/answers/i/87dae0f2a5d3452f951a3c0645e0a9e0_A.jpeg?a=answers&mr=0&x=1428339706&s=1eb2cb7e15bd82c17279270be4b29cd9 The line "starts" in (-8,-1) and "ends" in (0,4) So, [4-(-1)]/[0-(-8)] = [4+1]/[0+8]=5/8

OpenStudy (anonymous):

Lol I had -5/8 i was close

OpenStudy (anonymous):

What is the slope of the line whose equation is –48 = 2x − 8y?